# Performance Analysis of Indoor VLC Network with Secure Downlink NOMA for Body Blockage Model

### v3

@@ Abstract
In this work, we investigate the performance of indoor visible light communication (VLC) networks based on power domain non-orthogonal multiple access (NOMA) for mobile devices, where multiple legitimate users are equipped with photodiodes (PDs). We propose a body blockage model for both the legitimate users and eavesdropper to address scenarios where the communication links from transmitting light-emitting diodes (LEDs) to receiving devices are blocked by the bodies of all parties. Furthermore, in order to improve the secrecy without requiring knowledge of the channel state information (CSI) of the eavesdropper, we propose a novel LED arrangement that reduces the overlapping areas covered by different LED units supporting distinct users. We also suggest two LED transmission strategies, i.e., simple and smart LED linking, and compare their performance with the conventional broadcasting in terms of transmission sum rate and secrecy sum rate. Through computer simulations, we demonstrate the superiority of our proposed strategies to the conventional approach.
@@ I Introduction
Visible light communication (VLC) is a promising technology for the beyond fifth-generation (B5G) and sixth-generation (6G) networks [1]. Operating in the visible light spectrum, VLC utilizes the light emitted for both illumination and data transmission. With its high potential data rate, VLC has garnered significant interest [2]. Implemented through intensity modulation (IM) of light-emitting diodes (LEDs), VLC leverages existing infrastructure for illumination. It serves as an intriguing complement to radio frequency (RF) communication, particularly due to the need for strict RF radiation monitoring.
Non-orthogonal multiple access (NOMA) has gained attention for its high spectral efficiency compared to the conventional orthogonal multiple access (OMA) techniques [3, 4, 5]. Unlike OMA, NOMA allocates the same time-frequency resources to multiple users, maximizing the limited frequency spectrum utilization [6]. This work focuses on power domain NOMA, employing multi-user superposition coding at the transmitter and successive interference cancellation (SIC) at the receivers [7]. Power allocation algorithms for NOMA have been investigated in previous studies, e.g., [8, 9].
In addition, physical layer security (PLS) facilitates secure communication by leveraging wireless channel randomness, including fading, noise, and interference [10]. PLS aims to utilize the stochastic properties of wireless channels to provide secure communication. It ensures secrecy when the channel of the legitimate user surpasses that of the eavesdropper, making eavesdropping-resistant information transmission feasible without the need for symmetric keys or complex encryption/decryption algorithms [11, 12, 13, 14]. Also, PLS with VLC-enabled NOMA has been investigated in [15, 16, 17, 18]. Specifically, in [15], a precise expression of secure outage probability was derived for a NOMA-enabled multi-user multiple-input single-output (MISO) VLC system. In [16], secure transmission of MISO with NOMA under VLC, considering both known and unknown channel state information (CSI) of eavesdroppers, was investigated. In [17], secrecy performance of a NOMA-enabled VLC system in the presence of an active eavesdropper was demonstrated. In [18], secrecy performance is enhanced in the NOMA-enabled VLC system by the pseudo user, which interrupts the SIC process of the eavesdropper.
@@ I-A Literature Review and Motivation
In recent years, a variety of signal processing approaches has been investigated for VLC-aided PLS systems such as beamforming [19, 20, 21], precoding [22, 23], artificial optical jamming [24], and polar code-based secure coding [25]. Regarding mathematical tools, stochastic geometry theory is used to derive the average secrecy capacity of the system while the users are randomly distributed across the room [26]. Convex optimization theory can be used to enhance the secrecy performance of VLC systems that are subject to various illumination constraints, e.g., the LED optical power constraint [27], the peak power constraint to reduce eye damage [19], and the light energy harvesting and dimming control constraint [20]. However, most of the studies on PLS for VLC networks have focused on defending against external malicious eavesdroppers [23, 26, 27], some of which addressing the transmission of confidential information where users are unaware of information not intended for them [22]. Based on the properties of VLC channels, these studies considered applications of beamforming and jamming for enhancing the secrecy in the physical layer with high computational complexity. In practice, however, the position of the passive eavesdropper is unknown since they do not respond to transmission mechanisms. Furthermore, it is still highly complex for broadcasting systems to implement beamforming schemes in VLC transmission. With respect to LED arrangements, most works adopt a square lattice alignment to improve illuminance [28]. However, in broadcasting VLC, such an arrangement may degrade communication reliability since interference may occur in the overlapping areas served by multiple LED devices with each transmitting their own signal.
Furthermore, previous researches on indoor VLC systems ignore the effects of body blockage from users, with few exceptions such as [29, 30, 31]. In [29], the authors investigated the cylinder model of user bodies without considering the angles. Reference [30] investigates the use of reconfigurable intelligent surfaces (RIS) in order to avoid body blockage in VLC system. In [31], the modeling of body blockage has been investigated by dividing the orientation angle into several parts, each of which was assessed for blockage.
@@ I-B Contributions
  • In the aforementioned works (e.g., [31, 33]), the modeling of body blockage has only been partially addressed. However, the modeling methods of body blockage in these works are not universally applicable, i.e., these models are based on the statistical methods to approximate the channel state from LED sources to PD receivers. Therefore, we propose a geometrical-based general blockage model in the presence of a passive eavesdropper. This model accounts for scenarios where the light path from the LED to the PD may be blocked by user bodies, with blockage events determined through geometric calculations.
  • A new power allocation scheme is proposed based on NOMA to maximize the transmission sum rate of legitimate users, considering only their locations and statistical radiation angles. The system, however, is unaware of the devices’ azimuth and polar angles, which are important for detecting whether the light is blocked. Furthermore, two power allocation schemes are compared through computer simulations.
  • In order to decrease the overlapping area of different LED units, we propose a new LED arrangement, where LEDs are placed in the form of an equilateral triangle lattice. Furthermore, we introduce two transmission strategies for this LED arrangement, called simple and smart LED linking strategies, so as to investigate the trade-off between the transmission and secrecy sum rates.
@@ III-B Transmission and Secrecy Sum Rate
where the scaling factor $\frac{1}{2}$ is due to the baseband signal transmission and $\gamma_{k}(\theta_{k,n},\psi_{k,n},{\cal K}_{n})$ is the received signal-to-interference-plus-noise ratio (SINR) at the $k$th user (or eavesdropper) under the assumption of the perfect SIC, defined in Eq. 17 located at the top of the next page, with $|{\cal K}_{n}|$ representing the cardinality of the set ${\cal K}_{n}$. We define a new set for all the indices of LED light signal received by the $k$th user (or eavesdropper) in the set ${\cal K}_{n}$ as $\mathcal{I}_{k}=\{n\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in\mathcal{K}_{n},n\in\mathcal{I}\}$, e.g., if the $k$th user received the signals transmitted by LED sources with indices 1 and 2, then ${\cal I}_{k}=\{1,2\}$. Also, $\mathcal{I}_{k}^{\tilde{n}}$ is defined in the same manner as $\mathcal{I}_{k}$ with an additional condition $\tilde{n}\neq n$, i.e., $\mathcal{I}_{k}^{\tilde{n}}=\{n\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in\mathcal{K}_{n},n\in\mathcal{I},n\neq\tilde{n}\}$.
@@ IV-A3 Smart LED Linking Strategy
6:     ${{\mathbb{L}}}\leftarrow{{\mathbb{L}}}\cup\{{\cal L}_{n}\}$triangleright Generate a new set for the $n$th LED
@@
14:         if ${\cal L}_{l}\cap{\cal K}_{n}\neq\varnothing$ then $\triangleright$ Check any user indices duplicated between ${\cal S}_{l}$ and ${\cal S}_{n}$
15:              ${\cal M}\leftarrow{\cal M}\cup\{l\}$, ${\cal K}_{n}\leftarrow{\cal L}_{l}\cup{\cal K}_{n}$triangleright Record LED index to ${\cal M}$, and merge ${\cal L}_{l}$ to ${\cal K}_{n}$
@@ IV-B2 Maximum Sum Rate-based Scheme
where $\widehat{\gamma}_{k}\left(\overline{\lambda}_{k},\tilde{\omega}_{k,n},\beta_{k}^{\mathrm{opt}}({\cal K}_{n}),{\cal K}_{n}\right)$ is defined in Eq. 26 located at the top of this page

### v4

@@ Abstract
In this work, we investigate the performance of indoor visible light communication (VLC) networks based on power domain non-orthogonal multiple access (NOMA) for mobile devices, where multiple legitimate users are equipped with photodiodes (PDs). We propose a body blockage model for both the legitimate users and eavesdropper to address scenarios where the communication links from transmitting light-emitting diodes (LEDs) to receiving devices are blocked by the bodies of all parties. Furthermore, we propose a novel LED arrangement that improves secrecy without requiring knowledge of the channel state information (CSI) of the eavesdropper. This arrangement reduces the overlapping areas covered by different LED units supporting distinct users. We also suggest two LED transmission strategies, i.e., simple and smart LED linking, and compare their performance with the conventional broadcasting in terms of transmission sum rate and secrecy sum rate. Through computer simulations, we demonstrate the superiority of our proposed strategies to the conventional approach.
@@ I Introduction
Visible light communication (VLC) is a promising technology for the beyond fifth-generation (B5G) and sixth-generation (6G) networks [1]. VLC operates in the visible light spectrum, utilizing the emitted light for both illumination and data transmission. With its high potential data rate, VLC has garnered significant interest [2]. Implemented through intensity modulation (IM) of light-emitting diodes (LEDs), VLC leverages existing infrastructure for illumination. It serves as an intriguing complement to radio frequency (RF) communication, particularly due to the need for strict RF radiation monitoring.
Non-orthogonal multiple access (NOMA) has gained attention for its high spectral efficiency compared to the conventional orthogonal multiple access (OMA) techniques [3, 4, 5]. Unlike OMA, NOMA allocates the same time-frequency resources to multiple users, maximizing the limited frequency spectrum utilization [6]. This work focuses on power domain NOMA, employing multi-user superposition coding at the transmitter and successive interference cancellation (SIC) at the receivers [7]. Previous studies, such as [8, 9], have investigated power allocation algorithms for NOMA.
In addition, physical layer security (PLS) facilitates secure communication by leveraging wireless channel randomness, including fading, noise, and interference [10]. PLS aims to utilize the stochastic properties of wireless channels to provide secure communication. It ensures secrecy when the channel of the legitimate user surpasses that of the eavesdropper, making eavesdropping-resistant information transmission feasible without the need for symmetric keys or complex encryption/decryption algorithms [11, 12, 13, 14]. Also, PLS with VLC-enabled NOMA has been investigated in [15, 16, 17, 18]. Specifically, in [15], a precise expression of secure outage probability was derived for a NOMA-enabled multi-user multiple-input single-output (MISO) VLC system. In [16], the secure transmission of MISO with NOMA under VLC was investigated, taking into account both known and unknown channel state information (CSI) of eavesdroppers. In [17], secrecy performance of a NOMA-enabled VLC system in the presence of an active eavesdropper was demonstrated. It is shown in [18] that the pseudo user, who interrupts the eavesdropper’s SIC process, enhances secrecy performance in the NOMA-enabled VLC system.
@@ I-A Literature Review and Motivation
In recent years, a variety of signal processing approaches has been investigated for VLC-aided PLS systems such as beamforming [19, 20, 21], precoding [22, 23], artificial optical jamming [24], and polar code-based secure coding [25]. Stochastic geometry theory serves as a mathematical tool to determine the average secrecy capacity of the system, considering a random distribution of users throughout the room [26]. Convex optimization theory enhances the secrecy performance of VLC systems subject to various illumination constraints, such as the LED optical power constraint [27], the peak power constraint to reduce eye damage [19], and the light energy harvesting and dimming control constraint [20]. However, most of the studies on PLS for VLC networks have focused on defending against external malicious eavesdroppers [23, 26, 27], some of which addressing the transmission of confidential information where users are unaware of information not intended for them [22]. These studies, based on the properties of VLC channels, explored the use of beamforming and jamming techniques to enhance secrecy in the physical layer, despite the high computational complexity involved. In practice, however, the position of the passive eavesdropper is unknown since they do not respond to transmission mechanisms. Furthermore, broadcasting systems still face significant challenges in implementing beamforming schemes for VLC transmission. With respect to LED arrangements, most works adopt a square lattice alignment to improve illuminance [28]. However, in broadcasting VLC, such an arrangement may degrade communication reliability since interference may occur in the overlapping areas served by multiple LED devices with each transmitting their own signal.
Furthermore, previous researches on indoor VLC systems ignore the effects of body blockage from users, with few exceptions such as [29, 30, 31]. In [29], the authors investigated the cylinder model of user bodies without considering the angles. Reference [30] investigates the use of reconfigurable intelligent surfaces (RIS) in order to avoid body blockage in VLC system. The authors of [31] investigated body blockage modeling by dividing the orientation angle into several parts and assessing each part for blockage.
@@ I-B Contributions
  • The aforementioned works (e.g., [31, 33]) only partially address the modeling of body blockage. However, the modeling methods for body blockage in these works are not universally applicable, as they rely on statistical methods to approximate the channel state from LED sources to PD receivers. Therefore, we propose a geometrical-based general blockage model in the presence of a passive eavesdropper. It accounts for scenarios where user bodies may block the light path from the LED to the PD, determining blockage events through geometric calculations.
  • We propose a new power allocation scheme based on NOMA to maximize the transmission sum rate of legitimate users, taking into account only their locations and statistical radiation angles. We assume that the system is unaware of the azimuth and polar angles of the devices, crucial for determining if the light obstruction is present. Furthermore, two power allocation schemes are compared through computer simulations.
  • We propose a new LED arrangement that places LEDs in the form of an equilateral triangle lattice to reduce the overlapping area of different LED units. Furthermore, we introduce two transmission strategies for this LED arrangement, called simple and smart LED linking strategies, so as to investigate the trade-off between the transmission and secrecy sum rates.
@@ III-B Transmission and Secrecy Sum Rate
where the scaling factor $\frac{1}{2}$ is due to the baseband signal transmission and $\gamma_{k}(\theta_{k,n},\psi_{k,n},{\cal K}_{n})$ is the received signal-to-interference-plus-noise ratio (SINR) at the $k$th user (or eavesdropper) under the assumption of the perfect SIC, defined in Eq. 17 located at the top of this page, with $|{\cal K}_{n}|$ representing the cardinality of the set ${\cal K}_{n}$. We define a new set for all the indices of LED light signal received by the $k$th user (or eavesdropper) in the set ${\cal K}_{n}$ as $\mathcal{I}_{k}=\{n\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in\mathcal{K}_{n},n\in\mathcal{I}\}$, e.g., if the $k$th user received the signals transmitted by LED sources with indices 1 and 2, then ${\cal I}_{k}=\{1,2\}$. Also, $\mathcal{I}_{k}^{\tilde{n}}$ is defined in the same manner as $\mathcal{I}_{k}$ with an additional condition $\tilde{n}\neq n$, i.e., $\mathcal{I}_{k}^{\tilde{n}}=\{n\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in\mathcal{K}_{n},n\in\mathcal{I},n\neq\tilde{n}\}$.
@@ IV-A3 Smart LED Linking Strategy
6:     ${{\mathbb{L}}}\leftarrow{{\mathbb{L}}}\cup\{{\cal L}_{n}\}$
@@
14:         if ${\cal L}_{l}\cap{\cal K}_{n}\neq\varnothing$ then
15:              ${\cal M}\leftarrow{\cal M}\cup\{l\}$, ${\cal K}_{n}\leftarrow{\cal L}_{l}\cup{\cal K}_{n}$
@@ IV-B2 Maximum Sum Rate-based Scheme
where $\widehat{\gamma}_{k}\left(\overline{\lambda}_{k},\tilde{\omega}_{k,n},\beta_{k}^{\mathrm{opt}}({\cal K}_{n}),{\cal K}_{n}\right)$ is defined in Eq. 26 located at the top of the next page